OpenStudy (anonymous):

Rewrite the equation in Ax+By = C form. Use integers for A, B, and C. y-1 = 3(x+6)

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (skullpatrol):

Hint use the idea of balance to rearrange the terms in the equation: y-1 = 3(x+6) Ax+By = C

OpenStudy (skullpatrol):

But first use the distributive property: $$\Huge a(b+c)=ab+ac$$ $$\Huge 3(x+6)=?$$